Anton McClure

Location: Akron, OH, US

                     Web site: https://www.aperture.akron.oh.us/~asm/
                      E-Mails: asm@aperture.akron.oh.us, asm@gnu.org

Skills

Programming languages: C, VB.Net., C++, C#. Scripting languages: Perl, BASH, VBScript, PHP, SH, CSH, TCSH, KSH. Web servers: Aperture HTTPD, Apache (2.2 and 2.4), Internet Information Services (IIS), lighttpd, LightSpeed Web Server, thttpd, Abyss Web Server, httpi, NGINX. E-Mail servers: Sendmail, Exim 4, Courier, Postfix, Dovecot, Microsoft Exchange, Microsoft 365. DNS servers: ISC BIND 9 (NAMED). Web development (coded): HTML/HTML5, XHTML/XHTML5, Server-Side Includes, CSS/CSS3, CGIs, Custom CMS software. Web development (other): SEO, Google Analytics, Microsoft Clarity, Oracle AddThis. Document typesetting: SGML, XML, LaTeX. Databases: MySQL/MariaDB, SQLite, PostgreSQL, MongoDB, Microsoft SQL, Microsoft Access. Microsoft Windows: System administration and support skills for Windows servers and clients. GNU, BSD, & other UNIX: System administration and support skills with the GNU Operating System (Devuan, OpenMandriva Lx, Debian, RHEL / Oracle Linux / CentOS / Rocky Linux / AlmaLinux, Trisquel GNU/Linux, Slackware, Arch, Gentoo), FreeBSD, NetBSD, OpenBSD, and Sun/Oracle Solaris (including OpenSolaris, OmniOS CE, & OpenIndiana). Platforms & services: Oracle PeopleSoft, Microsoft Azure, Microsoft Intune/Endpoint/Entra, Microsoft 365 Security, University of Akron’s UAnet Services (Auth), OVHcloud, Hetzner Cloud. Software & tools: Microsoft Visual Studio, CVS , RCS, Git, Make, Microsoft 365, GNU GCC, Clang, OpenLDAP. Hardware: Repairing and replacing hardware, cleaning devices, maintaining RAID arrays, and re-imaging servers, clients, and other devices.

Work

Student Assistant, College of Business, University of Akron (August 2024-Present)

https://www.uakron.edu/cba/

• Collaborated with department chairs, administrative staff, and teaching faculty to identify and resolve a wide range of issues, ensuring academic operations while improving overall user experiences.

• Developed and implemented an efficient digital calendar signage solution, enabling faculty and students to access and engage with college events, improving academic efficiency and communication.

• Provided the Professional Development department with an improved method of managing and maintaining digital signage computers, improving efficiency and reliability for the department, the Labs student assistants, and the business students.

• Assisted in developing games for students at the Taylor Institute of Direct Marketing in their down time.

Lead Maintainer, GNU, FSF (February 2024-Present)

https://www.gnu.org/software/gnutrition/

• The Lead Maintainer of GNUtrition.

• Rewrote the program to bring GNUtrition to more people, remove the Python 2 and built-in GTK dependencies, and make the overall software more useful to more people.

• Wrote comprehensive documentation utilizing the GNU documentation standards.

Student Assistant, Information Technology Services, University of Akron (January 2023-May 2024)

https://www.uakron.edu/its/

• A student assistant at the University of Akron’s Information Technology Service Desk (ITSD). Started as an intern.

• Performed phone and remote technical support, computer repair, customer service, and other duties as assigned.

Projects

Aperture (2014-Present)

https://www.aperture.akron.oh.us/

• A collection of computers and computing services which runs DNS, web, e-mail, shell, and file servers.

• The Ocelot server, a private GNU (Devuan) shell server, initially started as an unnamed home-hosted shell system in 2014 that was powered by a HP laptop. In 2024, it was then moved to actual enterprise (Dell) hardware.

• Used directly with technologies at the University of Akron College of Business (See “Student Assistant, College of Business, University of Akron” under “Work”)

Flan Local-Access Notifier (2025-Present)

https://www.aperture.akron.oh.us/~asm/flan/

• Digital calendar signage software. Used at the College of Business.

• Designed to run as efficiently as possible.

Flan Local-Access Notifier Calendar Updater Program (2025-Present)

https://www.aperture.akron.oh.us/~asm/flan-cup/

• A utility to simplify Flan.

• Written for Microsoft Windows for use at the College of Business.

• Outputs to Flan XML and ICS file formats.

LibreCSS (2025-Present)

https://www.librecss.com/

• A collection of free web code and site designs.

• Uses as little client-side scripting as possible.

Education

College of Engineering and Polymer Science, University of Akron (May 2023-Present)

https://www.uakron.edu/computer/cis/

• A Bachelor’s of Science in Computer Information Systems for the programming degree option.

• Currently attending for this degree.

College of Engineering and Polymer Science, University of Akron (Jan 2019-May 2023)

https://www.uakron.edu/computer/cis/

• An Associate of Applied Business degree in Computer Information Systems for the programming degree option.

• Graduated with degree in May of 2023. Buchtel College of Arts and Sciences, University of Akron (Jan 2018-Jan 2019)

https://www.uakron.edu/computer/cs/

• A Bachelor’s of Science in Computer Science for the systems degree option.

• Transferred to associate degree for Computer Information Systems in December of 2018.

Notes

Links to latest version may be found on-line at https://www.aperture.akron.oh.us/~asm/cv.html.

Last updated on February 20, 2026.